OpenStudy (anonymous):

at a rodeo the total paid attendance was 3000. If the number of people who received a discount on the admission fee was 482 less than the number who did not, determine the number of paid attendees that paid in full.

OpenStudy (callisto):

let x be the no of attendees that paid in full From ''the number of people who received a discount on the admission fee was 482 less than the number who did not'' no of ppl who received discount = x-482 total no of ppl = 3000 no of attendees paid in full + no of attendees received discount = 3000 x + (x-482 ) =3000 2x = 3000+482 Can you solve it from here?

OpenStudy (anonymous):

@endiia Let us know if you need more help, will you!